SpaceRef

SpaceRef


NASA ISS On-Orbit Status 17 July 2013

Status Report From: NASA HQ
Posted: Wednesday, July 17, 2013

image New Status: Extravehicular Activity (EVA) Mobility Unit (EMU) S/N 3011 Troubleshooting: FE-5 Parmitano performed a subset of the Airlock deconfiguration procedure following yesterday's EVA by cleaning and stowing the prebreathe masks and returning the emergency books to their nominal location. He then began troubleshooting the EMU he wore yesterday during the EVA. Preliminary results of the troubleshooting tests indicate that water is likely being introduced into the oxygen ventilation circuit of the EMU. Today the crew performed a dry out of the EMU loop. Ground teams continue discussing forward troubleshooting plans to isolate the exact source of the water.

Post-EVA Cleanup: FE-3 Cassidy performed a nominal dump and fill of the feedwater tanks in the EMU he used during yesterday's EVA (S/N 3010).

Recycle Tank Remove and Replace (R&R): FE-6 Nyberg completed a nominal Recycle Tank R&R. The removed tank was then drained to a water container [ЕДВ] for stowage.

Payload Cleanup: Nyberg performed cleanup of the Advanced Biology Research System (ABRS) and Mice Drawer System (MDS) payload facilities in advance of upcoming payloads using these facilities.

Today's Planned Tasks

Daily Planning Conference (S-band) [Complete]
Gather two adjustable length hoses (П-М33-М32) for СРВ-К2М [Complete]
VIZIR. Experiment Ops. Tagup (S-band) [Complete]
Post-EVA Pille Dosimeter Reading [Complete]
Fungistat Treatment of Areas Behind FGB Panels / r/g 2906 [Complete]
ESA Weekly Crew Conference [Complete]
HAM Radio Power Up [Complete]
Replacement of adjustable length hose (П-М33-М32) for СРВ-К2М. [Complete]
ТКГ 418 (DC1) Stowage and IMS Ops [Complete]
SLAMMD [Complete]
EVA - Hatch Tool Relocate [Complete]
EMU-A/L - Post Extravehicular Activity (EVA) Deconfiguration [Complete]
СОЖ Maintenance
ТКГ 418 (DC1) Stowage and IMS Ops [Complete]
Fungistat Treatment of Areas Behind FGB Panels [Complete]
ALBEDO. Battery Charge (start) [Complete]
ИП-1 Flow Sensor Position Check [Complete]
Life On The Station Photo and Video [Complete]
WHC Fill [Complete]
RGN-REC TK - Transfer and Replacement [Complete]
IMS Update [Complete]
ALBEDO. Photo Spectral System (ФСС) Prep, Imagery, [Complete]
Replacement of ПФ1, ПФ2 Dust Filters and Cleaning В1, B2 Fan Screens in MRM2 [Complete]
В3 Fan Screen Cleaning in MRM2 [Complete]
EMU - Full Water Tank Dump and Fill [Complete]
RGN - Initiate drain into EDV [Complete]
RGN - Terminate drain into EDV [Complete]
Post-Exercise Hygiene Procedure [Complete]
Daily Planning Conference (S-band) [Complete]
Task List

CEO ops [In work]
WHC KTO Replace [Complete]
WHC EDV R&R
Roscosmos Report
Uragan Report
EKON Photo
Lab PCS HD Swap [Complete]
Regen WSTA Fill
Ground

Nominal Commanding
Three-Day Look Ahead:
Thursday, 7/18: EMU loop scrub, JSL 4.0 transition, ITCS sample gather/stow,SLAMMD
Friday, 7/19: 34S Emergency Drill, OGS H2 Sensor R&R/AAA Cleaning, EVA Tool Stow
Saturday, 7/20: Crew day off

QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - On
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 -Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up









// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks